# Structural characterizations and Obstructions to Planar-Rips complexes

### v1

@@ Abstract.
Given a scale parameter $r>0$, the Vietoris-Rips complex of $X\subset\mathbb{R}^{2}$ (referred to as planar-Rips complex) under the usual Euclidean metric is a (finite) simplicial complex whose simplices are subsets of $X$ with diameter at most $r$. This paper focuses on characterizing the simplicial complexes that can or cannot be realized as planar-Rips complexes. Building on the prior work of Adamszek et al., we classify, up to simplicial isomorphism, all $n$-dimensional pseudomanifolds and weak-pseudomanifolds that admit a planar-Rips structure, and further characterize two-dimensional, pure, and closed planar-Rips complexes. Additionally, the notion of obstructions to planar-Rips complexes has been introduced, laying the groundwork for algorithmic approaches to identifying forbidden planar-Rips structures. We also explore the correlations between planar-Rips complexes and the class of intersection graphs called disk graphs, establishing a natural isomorphism between the two classes. Parallely, our findings on planar-Rips complexes have been consolidated in terms of unit disk graphs for interested readers, depicting the significance of the topological and algebraic approaches. Several structural and geometric properties of planar-Rips complexes have been derived that are of independent interest.
@@
Vietoris-Rips complex, pseudomanifold, Euclidean plane, unit disk graph, hereditary, obstruction
@@ 1. Introduction
The Vietoris-Rips complex is a fundamental construct in topological data analysis (TDA), serving as a bridge between discrete data and continuous shape analysis. With an apt representation of certain networks or practical systems by a point set in Euclidean space, one can capture the intrinsic topological properties that are often obscured. TDA employs the Vietoris-Rips complex to infer these properties, encoding the data’s shape into a combinatorial framework. This in turn, facilitates the study of intricate geometry and connectivity within the system or the network at various scales [23].
This paper is specifically streamlined towards the study of Vietoris-Rips complexes of planar point sets. With the appealing interplay of geometry, topology, and combinatorics, this study is partly motivated by the ability of planar-Rips complexes to faithfully represent proximity-based relationships in the Euclidean plane, making them a useful tool for modeling various phenomena. To name a few, these complexes find their potential applications in wireless and sensor networks [13, 14, 15, 7], social network analysis [18], swarm robotics [17], and motion planning[16] where proximity and connectivity play pivotal roles in modeling interactions. Formally, the planar-Rips complex can be defined as follows:
Definition 1.1 (Planar-Rips complex).
Given a finite subset $X\subset\mathbb{R}^{2}$ with the usual Euclidean metric and a scale parameter $r>0$, the Vietoris-Rips complex of planar point set $X$ (also called planar-Rips complex) is a (finite) simplicial complex whose simplices are subsets of $X$ with diameter at most $r$. We shall denote it by $\mathcal{R}(X;r)$ [1].
One of the fundamental problems in TDA is the characterization of the planar-Rips complexes, at least up to homotopy. Our study is primarily driven by the works of [19, 20, 1, 2, 3] where the respective authors have widely addressed the characterization problem of planar-Rips complexes at various scales. In his paper [19], Hausmann demonstrated that up to homeomorphism, every $n$-sphere can be represented as a planar-Rips complex for some subset of circle $\mathbb{S}^{1}$. Adamszek and Adams in [3] obtain a wide range of homotopy types of planar-Rips complexes for various subsets of circle $\mathbb{S}^{1}$. In [1], Chambers et al. prove that the planar-Rips complexes have free fundamental group. As a consequence, one can verify using long exact sequences that every finite $2$-dimensional planar-Rips complex has homotopy type of wedge of circles and $2$-spheres. However, the problem of whether every planar-Rips complex is homotopy equivalent to wedge of spheres still remains open [2, Problem 7.3].
@@
A finite simplicial complex $K$ with the vertex set $V$ is said to admit a planar-Rips structure if there exists an injective map $f:V\to\mathbb{R}^{2}$ such that $K$ is isomorphic to the planar-Rips complex $\mathcal{R}(f(V);r)$ for some $r>0$.
This paper is centered around determining the simplicial complexes that do not admit planar-Rips structure. Therefore, with the scale-invariance under consideration, we fix $r=1$ throughout and use the notation $\mathcal{R}(X)$ instead of $\mathcal{R}(X;1)$.
Our paper builds on the prior work by M. Adamaszek et al. of classifying all normal pseudomanifolds admitting planar-Rips structure [2, Theorem 6.3]; They show that the normal psedomanifolds which admit planar-Rips structure are precisely the cross-polytopal spheres. In this paper, we extend this result to a larger class of pure simplicial complexes. Our main results are as follows:
@@
The n-dimensional cross-polytopal spheres ($n\geq 2$) are the only pseudomanifolds that admit planar-Rips structure.
This will be an inductive proof. We use this to further give a planar-Rips classification of $n$-dimensional weak-pseudomanifolds, whose each strongly connected component is a pseudomanifold.
@@
An n-dimensional weak-psedomanifold ($n\geq 2)$ admitting a planar-Rips structure is isomorphic to an iterated chain of $n$-dimensional cross-polytopal spheres. Further, they are homotopy equivalent to wedge sum of circles and $n$-spheres.
The tools from [1, 2] and the proximity relations derived towards proving the above two theorems are further applicable towards the characterization of $2$-dimensional pure and closed simplicial complexes that admit planar-Rips structure.
@@
Suppose $K$ is a $2$-dimensional pure and closed simplicial complex admitting a planar-Rips structure, then
  • (a)
@@
A significant implication of our study is towards characterization of a class of intersection graphs called unit disk graphs (UDGs). Notably, a natural isomorphism exists between the category of planar-Rips complexes and the category of unit disk graphs, with unit disk graphs being precisely the 1-skeleton of planar-Rips complexes. The literature on unit disk graphs and some of their applications can be found in [4, 5, 6]. In light of this isomorphism, the graph analogs to Theorems A,B,C have been substantiated in Section 5. For instance, the following theorem is a direct consequence of C, which significantly streamlines the recognition process for a class of unit disk graphs with all maximal cliques of size $3$.
Theorem D.
Let $G$ be a connected graph having clique number $\omega(G)=3$ and $\mathrm{degree}(e)\geq 2$ for every edge $e$ in $G$. If $G$ admits a unit disk graph structure, then
  • (a)
    $G$ contains an induced subgraph isomorphic to complete tripartite graph $K_{2,2,2}$.
  • (b)
    Every edge is contained in at most two distinct copies of $K_{2,2,2}$.
  • (c)
    Any edge $e$ with $\mathrm{degree}(e)\geq 3$ is contained in at least one copy of $K_{2,2,2}$.
@@
Another intriguing aspect of planar-Rips complexes is that they satisfy hereditary property. This means that every induced subcomplex is also a planar-Rips complex. It is well established that any hereditary class of simplicial complexes can be fully characterized in terms of minimal forbidden induced subcomplexes called obstructions. These obstructions represent critical subcomplex patterns that cannot appear as induced subcomplexes within any complex in the class. For example, the hereditary class of flag complexes can be defined as the simplicial complexes that do not contain the boundary complex of any $n$-simplex ($n>1$) as an induced subcomplex. In essence, any simplicial complex that does not admit a planar-Rips structure further contains a ‘minimal’ forbidden induced subcomplex. The significance of defining the class of planar-Rips complexes in terms of obstructions lies in its potential for algorithmic simplification in identifying planar-Rips complexes. This, in turn, offers more insight into the limitations and restrictions of associated systems or networks.
@@
Following Theorems A,B,C, we have indeed obtained a wide spectrum of simplicial complexes that do not admit a planar-Rips structure. This opens up the possibility of obtaining a large collection of obstructions to planar-Rips complexes by extracting each one as an induced subcomplex from the forbidden planar-Rips structures. The feasibility of establishing such a framework has been demonstrated by a few examples in Section 5. For example, Theorem 5.13 substantiates an obstruction to planar-Rips complex, derived from a triangulation of the projective space $\mathbb{R}P^{2}$.
In essence, the proximity relations and the structural/geometric characterizations established throughout this paper are of independent interest, laying the groundwork for an algorithmic approach to utilizing planar-Rips complexes.
@@
Organization of the paper: In Section 2, a background overview of planar-Rips complexes is presented by introducing key topological concepts and discussing the proximity relations for points in the Euclidean plane relevant to this paper. Section 3 delves into developing the necessary geometric aspects to prove A and B, leading to a complete classification of planar-Rips complexes with pseudomanifold and weak-pseudomanifold structures. Section 4 focuses on establishing the structural characterization of two-dimensional closed planar-Rips complexes as part of proving C. Lastly, in Section 5, we consolidate our findings on planar-Rips complexes in terms of unit disk graphs to prove D. In addition, the hereditary property of planar-Rips complexes has been explored and the notion of obstructions to planar-Rips complexes has been conceptualized. Illustrative examples are provided to demonstrate the utility of this concept.
@@ 2. Preliminaries
Throughout this paper, we consider the point set $X$ to be a finite subset of $\mathbb{R}^{2}$ and the planar-Rips complex $\mathcal{R}(X)$ to be connected. Further, the geometric realization of $\mathcal{R}(X)$ is considered while referring to the associated topological properties of $\mathcal{R}(X)$. We say two points/vertices are adjacent in $\mathcal{R}(X)$ if the distance between them is less than $1$ i.e. there is an edge between them in $\mathcal{R}(X)$.
@@
We now define the concept of “shadow” of a planar-Rips complex $\mathcal{R}(X)$ inside $\mathbb{R}^{2}$. This notion was originally introduced in a broader context for $\mathbb{R}^{n}$ by Chambers et al.[1] to simplify the computability. Likewise, we employ the concept of shadow to simplify the representation of proximity relations and certain topological properties of planar-Rips complexes.
Definition 2.1 (Shadow, projection map).
Given a finite set $X\subset\mathbb{R}^{2}$, the shadow $\mathcal{S}(X)$ is the image of the planar-Rips complex $\mathcal{R}(X)$ under the projection map $p:\mathcal{R}(X)\to\mathbb{R}^{2}$ which maps each vertex in $X$ to the corresponding point inside $\mathbb{R}^{2}$, and extends linearly to the simplices of $\mathcal{R}(X)$. More precisely,
$\mathcal{S}(X):=\bigcup_{\begin{subarray}{c}Y\subseteq X,\ |Y|\leq n+1\\ \mathrm{diam}(Y)<1\end{subarray}}\mathrm{conv}\ Y$
@@
Figure 1. An illustration of projection of planar-Rips complex $\mathcal{R}(X)$ onto its shadow $\mathcal{S}(X)$ inside $\mathbb{R}^{2}$
@@
[1] Let $X\subset\mathbb{R}^{2}$ be a finite subset and $p:\mathcal{R}(X)\to\mathcal{S}(X)$ be the projection map, then $p$ induces a $\pi_{1}$-isomorphism, hence implying that the fundamental group is free.
@@
[1, Proposition. 2.1] Suppose $X\subset\mathbb{R}^{2}$ and let $A,B,C,D$ be four points in $X$ such that the point $A$ is adjacent to $B$, and $C$ is adjacent to $D$ but not $A$ in $\mathcal{R}(X)$. Further, the segments $\overline{AB}$ and $\overline{CD}$ intersect in the shadow $\mathcal{S}(X)$. Then $B$ is adjacent to $D$ in $\mathcal{R}(X)$.
@@
The above result is a direct consequence of [4, Lemma 3.2], and has been elaborately explained in Section 5.
@@
Given $X\subset\mathbb{R}^{2}$, let $A$ be a subset of $X$ and $\sigma$ be a simplex such that $A\subseteq Lk(\sigma)$ inside $\mathcal{R}(X)$. Suppose a simplex $\gamma$ is in the convex hull of $A$, then $\gamma\in Lk(\sigma)$ and every vertex of $\gamma$ is adjacent to at least one vertex of $A$ in $\mathcal{R}(X)$.
@@
Let $u$ and $v$ be any vertex in $\sigma$ and $\gamma$ respectively. Firstly, by Caratheodory’s theorem, there exists vertices $a_{1},a_{2}\in A$ (not necessarily distinct) such that $v\in conv(u,a_{1},a_{2})$. Now, since $d(u,a_{i})\leq 1$ for all $i=1,2$, we have $d(u,w)\leq 1$ for any $w\in conv(u,a_{1},a_{2})$. Since this is true for any $u\in\sigma$, we have $\gamma\in Lk(\sigma)$.
@@
The above result generally holds even for the Vietoris-Rips complex of $X$ when $X\subset\mathbb{R}^{n}$ under the usual metric, and a similar proof is applicable.
@@
Suppose $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are any two components of link of a vertex inside $\mathcal{R}(X)$. By Lemma 2.6, one of the convex hulls of $V(C_{1})$ and $V(C_{2})$ cannot contain in the other. Further, from Lemma 2.4, the convex hulls cannot intersect on their boundaries. ∎

### v2

# On the Classification of Planar-Rips complexes and their corresponding unit disk graphs

@@ Abstract.
Given a metric space $(X,d)$, the Vietoris-Rips complex of $X$ at a scale of $r>0$ is a simplicial complex whose simplices are all those finite subsets of $X$ with diameter less than $r$. In this paper, we classify, up to simplicial isomorphism, all $n$-dimensional pseudomanifolds and weak-pseudomanifolds that can be realized as a Vietoris-Rips complex of planar point sets. We further classify two-dimensional, pure, and closed planar-Rips complexes up to homotopy. Additionally, we explore the hereditary properties and introduce the notion of obstructions in planar-Rips complexes. We also consolidate our findings to describe a class of unit disk graphs, having all maximal cliques of same cardinality. Several structural and geometric properties of planar-Rips complexes have also been derived.
@@
Vietoris-Rips complex, obstruction, unit disk graph, pseudomanifold, minimal cycle, closed
@@ 1. Introduction
Given a metric space $(X,d)$ and a scale parameter $\delta>0$, the Vietoris-Rips complex $\mathcal{R}(X;\delta)$ is a simplicial complex whose simplices are all those finite subsets of $X$ with diameter less than $\delta$, i.e., $\mathcal{R}(X;\delta)=\{A\subseteq X\colon|A|<\infty\text{ and }d(x,y)<\delta\ \forall x,y\in A\}$.The Vietoris-Rips complex is a fundamental construct in computational topology, serving as a bridge between discrete data and continuous shape analysis. First introduced by Vietoris [24] in the context of his work on (co)homology theory and covering spaces, and later independently by Rips in the context of geometric group theory [16], both used it essentially to understand complex topological spaces through a combinatorial framework. In this paper, we will only be working within the Euclidean setting, and all our points sets will be finite subsets of $\mathbb{R}^{2}$. Henceforth, we shall refer to them as planar-Rips complexes [13].
@@
A finite simplicial complex $K$ with the vertex set $V$ is said to admit a planar-Rips structure if there exists an injective map $f:V\to\mathbb{R}^{2}$ such that $K$ is isomorphic to the Vietoris-Rips complex $\mathcal{R}(f(V);\delta)$ for some $\delta>0$ under the usual Euclidean metric.
This paper is specifically streamlined towards determining the simplicial complexes that (do not) admit a planar-Rips structure. Therefore, with the scale-invariance under consideration, we fix $\delta=1$ throughout and use the notation $\mathcal{R}(X)$ instead of $\mathcal{R}(X;1)$. Our study is primarily driven by the works of [18, 19, 13, 3, 1] where the respective authors have widely addressed the characterization problem of planar-Rips complexes at various scales.
In his paper [18], Hausmann demonstrated that up to homeomorphism, every $n$-sphere admits a planar-Rips structure for some subset of circle $\mathbb{S}^{1}$. This further implies that for a given wedge of spheres, there exists a subset $X\subset\mathbb{R}^{2}$ whose planar-Rips complex is homotopy equivalent to the given wedge. In [13], Chambers et al. prove that the planar-Rips complexes have free fundamental group. As a consequence, one can easily verify that every finite two-dimensional planar-Rips complex is homotopy equivalent to wedge of circles and $2$-spheres.
In [1], Adamaszek and Adams obtain a wide range of homotopy types of planar-Rips complexes for subsets of circle $\mathbb{S}^{1}$. In particular, Adamaszek et al. in [2] show that the planar-Rips complex of any finite subset of $\mathbb{S}^{1}$ is homotopy equivalent to either a point, an odd sphere, or a wedge sum of spheres of the same even dimension. Further in [3], Adamaszek et al. show that the cross-polytopal spheres are the only normal pseudomanifolds that admit planar-Rips structure. With this, they postulate the following conjecture, a problem which still remains open:
Conjecture 1.2.
[3, Problem 7.3] For any $X\subset\mathbb{R}^{2}$, the planar-Rips complex $\mathcal{R}(X)$ is either contractible or homotopy equivalent to wedge of spheres.
In this spirit, we extend the planar-Rips classification to a larger class of pure simplicial complexes, thereby providing an affirmative evidence supporting the conjecture. Our main results are as follows:
@@
(See Theorem 3.7 &\& 3.15) The n-dimensional cross-polytopal spheres ($n\geq 2$) are the only pseudomanifolds that admit planar-Rips structure.
This will be an inductive proof. We use this to further give a planar-Rips classification of $n$-dimensional weak-pseudomanifolds, each of whose strongly connected component is a pseudomanifold. (See Section 3 for definition of pseudomanifold and weak-pseudomanifold)
@@
(See Theorem 3.18 &\& 3.19) An n-dimensional weak pseudomanifold ($n\geq 2)$ admitting a planar-Rips structure is isomorphic to an iterated chain of $n$-dimensional cross-polytopal spheres. Further, they are homotopy equivalent to wedge sum of circles and $n$-spheres.
The tools from [13, 3] and the proximity relations derived towards proving the above two theorems are further applicable towards the characterization of two-dimensional pure and closed simplicial complexes that admit planar-Rips structure.
@@
(See Theorem 4.7 &\& Proposition 4.14) Suppose $K$ is a two-dimensional pure and closed simplicial complex admitting a planar-Rips structure, then
  1. (a)
@@
As a simple consequence of the above theorem, if $K$ is a minimal $2$-cycle, then it isomorphic to octahedron. Furthermore, $K$ is homotopy equivalent to wedge of circles and $2$-spheres.
The planar-Rips complexes are also useful from an application standpoint because of their ability to faithfully represent proximity-based relationships in the Euclidean plane. Recall that a (unit) disk graph is the intersection graph of (unit) equal-radius disks in the Euclidean plane. Notably, a natural isomorphism exists between the category of planar-Rips complexes and the category of disk graphs, with disk graphs being precisely the 1-skeleton of planar-Rips complexes. For more on (unit) disk graphs, we refer the reader to [20, 14, 10, 21]
In light of the isomorphism, we use our findings to characterize a class of disk graphs with a uniform clique structure. In particular, the graph analogs to Theorems A,B,C have been substantiated in Section 5, which significantly streamline the graph recognition process for disk graphs.
@@
Another intriguing aspect of planar-Rips complexes is that they satisfy hereditary property. This means that every induced subcomplex is also a planar-Rips complex. It is well established that any hereditary class of simplicial complexes can be fully characterized in terms of minimal forbidden induced subcomplexes called obstructions. These obstructions represent critical subcomplex patterns that cannot appear as induced subcomplexes within any complex in the class. For example, the hereditary class of flag complexes can also be defined as the simplicial complexes that do not contain the boundary complex of any $n$-simplex ($n>1$) as an induced subcomplex. The study of obstructions and hereditary properties are found in [25, 28, 17, 27].
@@
In essence, any simplicial complex that does not admit a planar-Rips structure further contains a ‘minimal’ forbidden induced subcomplex. Following Theorems A,B,C, we have indeed obtained a wide spectrum of simplicial complexes that do not admit a planar-Rips structure. This opens up the possibility of obtaining a large collection of obstructions of planar-Rips complexes by extracting each one as an induced subcomplex from the forbidden planar-Rips structures. The feasibility of establishing such a framework has been demonstrated by a few examples in Section 5. For example, Theorem 5.12 substantiates an obstruction of planar-Rips complex, derived from a minimal flag triangulation of the projective space $\mathbb{R}P^{2}$.
Acknowledgments
We thank Henry Adams and Jan Segert for making the Mathematica Demonstration on Vietoris-Rips and $\check{C}ech$ complexes publicly available [4], which aided in refining our approaches. We also thank Priyavrat Deshpande and Anurag Singh for introducing the notion of obstructions to shellability during the NCMW workshop “Cohen-Macaulay Simplicial Complexes in Graph Theory”, which inspired the extension of this concept for planar-Rips complexes.
1.1. Organization of the paper
@@
In Section 2, a background overview and key topological concepts are introduced relevant to this paper. Section 3 delves into developing the necessary geometric aspects to prove Theorem A and Theorem B. Section 4 focuses on establishing the structural characterization of two-dimensional closed planar-Rips complexes as part of proving Theorem C. Lastly, in Section 5, we consolidate our findings on planar-Rips complexes in terms of unit disk graphs. In addition, the hereditary property of planar-Rips complexes has been explored and the notion of obstructions of planar-Rips complexes has been conceptualized with few illustrative examples.
@@ 2. Preliminaries
Throughout this paper, we consider the point set $X$ to be a finite subset of $\mathbb{R}^{2}$ and the planar-Rips complex $\mathcal{R}(X)$ to be connected. Further, the geometric realization of $\mathcal{R}(X)$ is considered while referring to the associated topological properties of $\mathcal{R}(X)$. We say two points/vertices are adjacent in $\mathcal{R}(X)$ if the distance between them is less than $1$ i.e., there is an edge between them in $\mathcal{R}(X)$.
@@
We now define the concept of “shadow” of a planar-Rips complex $\mathcal{R}(X)$ inside $\mathbb{R}^{2}$. This notion was originally introduced by Chambers et al.[13] to simplify the computability. Likewise, we employ the concept of shadow to simplify the representation of proximity relations and certain topological properties of planar-Rips complexes.
Definition 2.1 (Shadow, projection map [3]).
Given a finite set $X\subset\mathbb{R}^{2}$, the shadow $\mathcal{S}(X)$ is the image of the planar-Rips complex $\mathcal{R}(X)$ under the projection map $p\colon\mathcal{R}(X)\to\mathbb{R}^{2}$ which maps each vertex in $X$ to the corresponding point inside $\mathbb{R}^{2}$, and extends linearly to the simplices of $\mathcal{R}(X)$. More precisely,
$\mathcal{S}(X)\coloneq\bigcup_{\begin{subarray}{c}Y\subseteq X,\ |Y|\leq 3\\ \mathrm{diam}(Y)<1\end{subarray}}\mathrm{conv}\ Y$
Note that by Caratheodory’s theorem, the convex hull of any finite set $Y$ in $\mathbb{R}^{2}$ is equal to the union of convex hulls of at most 3-element subsets of $Y$. Henceforth, the condition $|Y|<\infty$ is redundant to $|Y|\leq 3$ in the above defintion.
@@
Figure 1. An illustration of projection of a planar-Rips complex $\mathcal{R}(X)$ onto its shadow $\mathcal{S}(X)$ inside Euclidean plane.
@@
[13] For any finite subset $X\subset\mathbb{R}^{2}$, the projection map $p\colon\mathcal{R}(X)\to\mathcal{S}(X)$ induces a $\pi_{1}$-isomorphism, hence implying that the fundamental group $\pi_{1}(\mathcal{R}(X))$ is free.
@@
[13, Proposition. 2.1] Suppose $X\subset\mathbb{R}^{2}$ and let $A,B,C,D$ be four distinct points in $X$ such that the point $A$ is adjacent to $B$, and $C$ is adjacent to $D$ but not $A$ in $\mathcal{R}(X)$. Further, the segments $\overline{AB}$ and $\overline{CD}$ intersect in the shadow $\mathcal{S}(X)$. Then $B$ is adjacent to $D$ in $\mathcal{R}(X)$.
@@
The above result is a direct consequence of [20, Lemma 3.2], and has been elaborately explained in Section 5.
@@
Given $X\subset\mathbb{R}^{2}$, let $A$ be a subset of $X$ and $\sigma$ be a simplex such that $A\subseteq\mathrm{Lk}(\sigma)$ inside $\mathcal{R}(X)$. Suppose a simplex $\gamma$ is in the convex hull of $A$, then $\gamma\in\mathrm{Lk}(\sigma)$ and every vertex of $\gamma$ is adjacent to at least one vertex of $A$ in $\mathcal{R}(X)$.
@@
Let $u$ and $v$ be any vertex in $\sigma$ and $\gamma$ respectively. Firstly, by Caratheodory’s theorem, there exists vertices $a_{1},a_{2}\in A$ (not necessarily distinct) such that $v\in conv(u,a_{1},a_{2})$. Now, since $d(u,a_{i})\leq 1$ for all $i=1,2$, we have $d(u,w)\leq 1$ for any $w\in conv(u,a_{1},a_{2})$. Since this is true for any $u\in\sigma$, we have $\gamma\in\mathrm{Lk}(\sigma)$.
@@
The above result generally holds even for the case when $X\subset\mathbb{R}^{n}$ under the usual Euclidean metric, and a similar proof is applicable.
@@
Suppose $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are any two components of link of a vertex inside $\mathcal{R}(X)$. By Lemma 2.6, one of the convex hulls of $V(\mathcal{C}_{1})$ and $V(\mathcal{C}_{2})$ cannot contain in the other. Further, from Lemma 2.4, the convex hulls cannot intersect on their boundaries. ∎